OpenStudy (anonymous):

what is the solution to In(a+3)=In(a-3)=In16

OpenStudy (anonymous):

Is it +/- in the middle of those "ln" operators?

OpenStudy (anonymous):

if it is + in the middle then, ln((a+3)(a-3))=ln16 a^2-9=16 a^2=25 a= +5 or - 5

OpenStudy (anonymous):

a can't be -5 because then ln(a-3) will be undefined.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@GT , you are right. Sorry I am liitle sleepy :P

OpenStudy (anonymous):

also if 3^2x=100 then to the nearest tenth what does x =

OpenStudy (anonymous):

aply log on both sides and then apply power rule

OpenStudy (anonymous):

how would you do that?

OpenStudy (anonymous):

log ( 3^(2x))= log (10^2) POwer rule is \[\log_{a}b^c = c \log_{a}b\] therefore (2x)log 3 = 2 log10 x= log 10 / log 3

OpenStudy (anonymous):

Thank you again

OpenStudy (anonymous):

Welcome :)

OpenStudy (anonymous):

the amount of a 10 miligram dose of drug remaining in the body is given by the formula 10(.8)^t. Find the nearest 10th of an hour how long it takes for half of the drug dose to leave the body

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

10(0.8)^t = 0.5 (0.8)^t= 0.05 Applying log on both sides t*log(0.8)=log(0.05) t= log(5*10^(-2))/log (8 * 10^-1) t=(log(5)-2log10) / (log(8) - 1 *log(10) t=(log (5) - 2) / ( log (8) - 1)

OpenStudy (anonymous):

@rama1010 , next time ask separately . Not in this thread :)

OpenStudy (anonymous):

ok but what would be the answer in decimal form

OpenStudy (anonymous):

Use calculator please and find the log values :). Substitute and simplify

OpenStudy (anonymous):

thank you very much for all your help. I am very grateful!!!

OpenStudy (anonymous):

however i still can't seem to find the actual values

OpenStudy (anonymous):

sorry? log 5 =0.698970004 log 8 = 0.903089987 Now substitute and simplify :)

OpenStudy (anonymous):

the answer is .777 right